INTENSITY CALCULATIONS FOR THE Ã^{1}A_{u}(C_{2h}) − X̃^{1}Σ_{g}^{+}(D_{∞h}) TRANSITION OF ACETYLENE

\underline{JAMES K. G. WATSON}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario, Canada K1A 0R6.

The Ã − X̃ transition of acetylene was the first electronic transition for which a change of point group was confirmed in detail^{a}. Previously, harmonic Franck-Condon calculations for this transition were presented^{b}. The present calculations allow for anharmonicity in the ground state, although the upper-state potential is still assumed to be harmonic^{c}. The transition moment is taken to be proportional to the q′′_{4} bending coordinate. For the calculation of the dispersed fluorescence spectrum^{d} it is found that the global potential surface of Halonen, Child and Carter^{e} gives the best qualitative agreement for the intensities, but could be improved for the energies of high bending states.

aC. K. Ingold and G. W. King, J. Chem. Soc., 2702–2755 (1953); K. K. Innes, J. Chem. Phys. 22, 863–876 (1954).

bJ. K. G. Watson, Paper TG11, OSU Symposium (1998).

cJ. D. Tobiason, A. L. Utz, E. L. Sibert III, and F. F. Crim, J. Chem. Phys. 99, 5762–5767 (1993).

dM. P. Jacobson, Ph.D. Thesis, Massachusetts Institute of Technology (1999); M. P. Jacobson and R. W. Field, J. Phys. Chem.

104, 3073 (2000).

eL. Halonen, M. S. Child, and S. Carter, Mol. Phys. 47, 1097 1982.